A2B4(g) - 2AB2(g) initial pressure of A2B4 in closed system was 0.85 atm and the final equilibrium pressure became 0.98 atm .so calculate Kp and Kc


1
Expert's answer
2021-01-04T03:47:43-0500

Solution:

1) Given chemical reaction:


A2B4(g) - 2AB2(g)

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 I & 0.85 & 0 \\ \hline
 C & -x & 2x \\
 \hdashline
 E & 0.98 & -0.26
\end{array}"


0.98-(-x)=0.85

x=- 0.13


"K_p=\tfrac{p[AB_2]^2}{p[A_2B_4]}=\tfrac{0.26^2}{0.98}"


"K_p=0.069"


2) Concentration constant:

In this case, the temperature is not given in this question, so we consider that we are in standard condition.


"K_p=K_c(RT)^{n(gas)}"


"n_{gas}=2-1=1"


"K_c=\tfrac{K_p}{RT}=\tfrac{0.069}{0.0821*273}"


"K_c=0.003"